# Distributed Quantum Magnetic Sensing for Infrastructure-free Geo-localization

### v1

@@ I Introduction
Robust and reliable navigation is a cornerstone of modern technology, supporting applications ranging from autonomous transportation to defense systems [nahavandi2025comprehensive, olthuis2025cyberattacks]. For decades, the Global Positioning System (GPS) and similar satellite constellations have served as the primary sources of global positioning. However, the reliance on these weak signals beamed from space introduces critical vulnerabilities. They are susceptible to jamming, spoofing, and are simply unavailable in environments such as underwater, underground, or in large buildings [zidan2020gnss, khan2021gps, radovs2024recent, tanaka2023first]. This fragility necessitates the development of complementary navigation technologies that remain robust in contested environments.
@@
One promising approach for navigation and localization is to use the Earth’s magnetic field [lei2025review, karshakov2020promising]. The Earth’s core, which is composed primarily of molten iron, acts as a colossal dynamo that produces our planet’s magnetic field. This interplay of iron and motion deep within the Earth generates magnetic fields strong enough to influence the surface, commonly referred to as the Earth’s main (core) magnetic field. Superimposed on this main field are magnetic anomalies. These irregularities are often caused by different geological formations and local variations in crustal magnetization. One significant contribution comes from magnetic minerals embedded within the Earth’s crust, which create localized disruptions that reveal hidden geological structures. While the earth’s magnetic field entails valuable geo-spatial structure, which many marine animals (e.g, sea turtles, whales) and migratory birds use for migration [packmor2024migratory, keller2021map, wiltschko2022discovery], the ultra-weak and subtle variations of the Earth’s core and anomaly magnetic field present unique challenges for civilian and military usage. With the development of highly sensitive quantum sensors, using the Earth’s magnetic field, specifically the magnetic anomalies, as a passive, globally available reference for navigation and localization may become a viable technical solution.
@@
Realizing this potential requires magnetometers with sufficient sensitivity and stability to resolve weak crustal signatures on top of a much larger background field [wang2022quantum, lonard2025limits, barry2020sensitivity]. The nitrogen–vacancy (NV) center in a diamond is an excellent candidate, formed by replacing one carbon atom in the lattice with a nitrogen atom and creating an adjacent vacant lattice site [li2025review]. The negatively charged NV-center is sensitive to the ambient magnetic field. It features an electronic spin-triplet ground state with three sub-levels labeled by the spin projection $m_{S}=\{0,\pm 1\}$, which can be initialized and read out optically. In a typical setup for magnetic sensing, the spin is first prepared in the $m_{S}=0$ state and then driven to coherent transitions (or superposition) between $m_{S}=0$ and $m_{S}=\pm 1$. An external magnetic field shifts the corresponding resonance frequencies (or the probability amplitudes between $m_{S}$’s) via the Zeeman effect. By monitoring the frequency, one can estimate the strength of the magnetic field with high precision [kuwahata2020magnetometer, hong2013nanoscale, abe2018tutorial]. This setup is a standard Ramsey interferometry for a single shot measurement [degen2017quantum]. Like classical sensing paradigms, $N$ independent executions of NV-center Ramsey measurements only lead to $\frac{1}{\sqrt{N}}$ error scaling due to the independent per-shot Gaussian noises. This limit is widely known as the standard quantum limit (SQL). In contrast, some recent works have demonstrated that NV-center quantum magnetometers with only correlated measurements (without using entanglement) can beat SQL and approach the Heisenberg limit (HL) with a scaling factor of $\frac{1}{N}$ [guo2025two, bonato2016optimized], which is the ultimate precision limit only constrained by the law of physics. Therefore, it enables highly sensitive sensing of the Earth’s magnetic field and results in accurate geo-localization.
@@ II Related Work
The foundation of any magnetic-based localization system is the magnetic anomaly map, which provides the reference field for positioning. Gupta et al. [gupta2024lower] established the relationship between magnetic anomaly map quality and achievable positioning accuracy via the CRLB. To implement the bound, they considered two measurement noise models: unimodal Gaussian noise and multimodal noise. Through simulations over a $250\times 400~\text{km}^{2}$ region, their results demonstrated a strong correspondence between higher magnetic gradients and lower CRLB (i.e., better accuracy). Recognizing that localization performance depends on the spatial variability of the magnetic field, Penumarti et al. [penumarti2024global] addressed localization uncertainty in magnetic anomaly navigation by introducing a global, uncertainty-aware path planner. Their key contribution was the generation of an entropy map from pre-existing magnetic anomaly map to statistically quantify areas of high information gain. They employed a multi-objective potential field planner that attracted the vehicle toward these low-entropy regions while simultaneously guiding it to a global goal. Hardware experiments using TurtleBot 4 equipped with a Quspin QTFM total-field magnetometer demonstrated that this method produced paths that significantly improved localization stability and reduced estimation covariance compared to existing methods.
@@
While map quality is important, accurate measurement of weak magnetic anomalies requires highly sensitive magnetometers. Bonato et al. [bonato2016optimized] demonstrated DC magnetic field sensing with single-spin NV-center magnetometer using a real-time adaptive measurement protocol. They implemented an adaptive Ramsey interferometry protocol in which the spin readout basis was dynamically updated after each single-shot measurement using a Bayesian estimator. The protocol achieved a remarkable magnetic field sensitivity of $6.1\pm 1.7~\text{nT}/\sqrt{\text{Hz}}$ over a dynamic range of $1.78~\text{mT}$. Building on the development of high-sensitivity NV-center magnetometers, Wang et al. [wang2022quantum] investigated their viability as aiding sensors for INS using total magnetic intensity maps for localization. However, in practical platforms, the weak crustal magnetic signal is often corrupted by vehicle-induced interference, making signal separation a critical challenge. To improve magnetic signal fidelity, the United States Air Force Research Laboratory, in collaboration with MIT Lincoln Lab, introduced the MagNav signal enhancement challenge [gnadt2020signal], which aimed to develop algorithms to separate the Earth’s magnetic anomaly field from strong aircraft-induced magnetic interference. A team from SandboxAQ addressed the challenge by proposing a physics-informed machine learning model that integrated the classical Tolles-Lawson (T-L) model with a Liquid Time-Constant Network (LTC) [nerrise2024physics]. Their approach used a closed-form continuous-time (CfC) variant of the LTC to model the nonlinear dynamics of the aircraft-induced magnetic noise. Experimental results demonstrated that the proposed LTC-CfC network achieved up to a $64\%$ reduction in compensation error compared to the standard T-L model and outperformed other baseline methods.
@@
Building on these foundations, several integrated systems have demonstrated practical MagNav capabilities. Q-CTRL introduced a full-stack system [muradoglu2025quantum] integrating a quantum scalar optically pumped atomic magnetometer, a classical fluxgate vector magnetometer, and a software stack for magnetic denoising and map matching. The software stack processes the magnetometer readings to perform real-time platform denoising and provides periodic position fixes by matching the cleaned magnetic signal against a known anomaly map (e.g., EMAG2v3). Across flight trials, this quantum-assured system consistently outperformed a strategic-grade INS, achieving superior bounded positioning accuracy with a best-case result of 22 meters. In ground trials, the system demonstrated a bounded positioning error approximately seven times lower than that of the INS. However, a major limitation of such map-based approaches is their dependence on pre-surveyed maps. To address this, Lee et al. [lee2020magslam] introduced MagSLAM, an aerial simultaneous localization and mapping framework (SLAM) that exploits the Earth’s magnetic anomaly field without relying on pre-surveyed magnetic maps. The authors adapted the FastSLAM algorithm (Rao-Blackwellized particle filter) [montemerlo2002fastslam] to fuse scalar magnetic anomaly measurements with an INS. Their method uses a nine-state INS error model to constrain drift, while simultaneously estimating the vehicle’s trajectory and building a local magnetic map without any prior survey. Flight tests were performed with a Cessna 208B equipped with a cesium-vapor scalar magnetometer, barometer, and high-grade INS over a $9\times 12~\text{km}$ region in Louisa, Virginia, at $150~\text{m}$ above ground level. The MagSLAM filter achieved an approximate $17~\text{m}$ distance root mean square position accuracy over a 100-minute flight.
@@ III Problem formulation
Figure 1: The CRLBs for a single-spin NV-center magnetometer and two commercial sensors (HMC5983 and AKM8975) [dammann2024cramer], with respect to sensing time per measurement shot $\tau$. The red curve shows the NV-center CRLB after a fixed total sensing time $T_{\mathrm{total}}=1.5~\mathrm{s}$ with $T_{2}^{\ast}=1500~\mu\mathrm{s}$ and readout fidelities $F_{0}=0.88$ and $F_{1}=0.98$.
@@ IV-B Quantum Cramér–Rao Lower Bound
To quantify the gained knowledge from this measurement, we use the Fisher information. The Fisher information measures how the outcome probabilities respond to changes in the parameter of interest, $\omega$. In our case, the Fisher information, and thus the measurement precision, depends on the controllable phase $\vartheta$ of the second $\pi/2$ pulse and is given as follows [nolan2021machine]:
@@ IV-C Classical Magnetometer Cramér–Rao Lower Bound
To demonstrate the quantum advantage of the NV-center magnetometer, we compare its CRLB against that of established classical magnetometers. Fig. 1 shows the quantum CRLB on magnetic field sensing as a function of sensing time $\tau$ per Ramsey shot, for a fixed total sensing time $T_{total}=1.5~\mathrm{s}$. In comparison, we set the inhomogeneous dephasing time $T_{2}^{*}=1,500~\mu\mathrm{s}$, consistent with room-temperature single-spin systems [herbschleb2019ultra]. The classical sensors used for this benchmark were selected from [dammann2024cramer], which are neither dominated by external magnetic field noise (e.g., MMC246xMT) nor exhibit anisotropic resolution (e.g., YAS532, YAS529). For each device, we use the noise variance reported in [dammann2024cramer] as the variance parameter in the classical model of Theorem 1, which yields the horizontal lines in Fig. 1. As shown in figure, the quantum CRLB is highly dependent on sensing time, achieving a minimum at an optimal value before increasing due to decoherence. The shaded regions highlight the operational windows where quantum sensor’s theoretical precision surpasses that of classical benchmarks, demonstrating quantum advantage. This trade-off motivates the need to operate the sensor in optimal sensing time.
@@ V-A System Overview
In this work, we deliberately adopt the same single-spin NV-center magnetometer setup as in our prior two-stage sensing protocol [guo2025two]. The rationale for selecting this protocol is that we have experimentally demonstrated that it can saturate the quantum CRLB as illustrated in Fig.1, thus beating SQL and approaching HL. More importantly, it achieves the best estimation accuracy among existing methods in the literature for estimating wide-range magnetic field.
@@ V-B Two-Stage Adaptive Sensing Protocol
We adopt the two-stage adaptive sensing protocol introduced in our earlier study [guo2025two] and use it as the front end of the magnetic localization pipeline. We refer readers to [guo2025two] for more details of the protocol while we only present a sketch of how it works here. The protocol comprises two phases: an offline training phase, followed by an online inference phase.
@@ VI-A Magnetic Map
In this work, we used the composite magnetic anomaly grid for the United States and Canada published by the USGS in March 2025 [mccafferty2023national]. This dataset merges magnetic surveys [ravat2010preliminary, ADGGS2016, MilesOneschuk2016] to create a magnetic anomaly map of North America with a resolution of approximately 1 km (30 arc-seconds). To extract the ROI from the broader North America dataset, we applied a geospatial clipping operation. A polygon defining the CONUS boundary was constructed from the Natural Earth Admin-1 states and provinces dataset [NaturalEarth2024]. This polygon was used to mask the original GeoTIFF, excluding Alaska, Hawaii, and Canadian territories. Fig. 3(a) shows the resulting ROI, spanning longitudes from approximately 125° to 66.5° and latitudes from 24° to 49.5°. The anomaly values within the CONUS ROI exhibit a wide dynamic range, from approximately 332.4-332.4 nT to over $430.4$ nT, reflecting the complex and varied geology of the continent.
@@ Appendix A Proof of Theorem 1: Classical Magnetometer Cramér–Rao Bound
The Fisher information $F(B)$ for a single parameter is defined as [dammann2024cramer]:

### v2

@@ I Introduction
Robust and reliable navigation is a cornerstone of modern technology, supporting applications ranging from autonomous transportation to defense systems [1, 2]. For decades, the Global Positioning System (GPS) and similar satellite constellations have served as the primary sources of global positioning. However, the reliance on these weak signals beamed from space introduces critical vulnerabilities. They are susceptible to jamming, spoofing, and are simply unavailable in environments such as underwater, underground, or in large buildings [3, 4, 5, 6]. This fragility necessitates the development of complementary navigation technologies that remain robust in contested environments.
@@
One promising approach for navigation and localization is to use the Earth’s magnetic field [7, 8]. The Earth’s core, which is composed primarily of molten iron, acts as a colossal dynamo that produces our planet’s magnetic field. This interplay of iron and motion deep within the Earth generates magnetic fields strong enough to influence the surface, commonly referred to as the Earth’s main (core) magnetic field. Superimposed on this main field are magnetic anomalies. These irregularities are often caused by different geological formations and local variations in crustal magnetization. One significant contribution comes from magnetic minerals embedded within the Earth’s crust, which create localized disruptions that reveal hidden geological structures. While the earth’s magnetic field entails valuable geo-spatial structure, which many marine animals (e.g, sea turtles, whales) and migratory birds use for migration [9, 10, 11], the ultra-weak and subtle variations of the Earth’s core and anomaly magnetic field present unique challenges for civilian and military usage. With the development of highly sensitive quantum sensors, using the Earth’s magnetic field, specifically the magnetic anomalies, as a passive, globally available reference for navigation and localization may become a viable technical solution.
@@
Realizing this potential requires magnetometers with sufficient sensitivity and stability to resolve weak crustal signatures on top of a much larger background field [12, 13, 14]. The nitrogen–vacancy (NV) center in a diamond is an excellent candidate, formed by replacing one carbon atom in the lattice with a nitrogen atom and creating an adjacent vacant lattice site [15]. The negatively charged NV-center is sensitive to the ambient magnetic field. It features an electronic spin-triplet ground state with three sub-levels labeled by the spin projection $m_{S}=\{0,\pm 1\}$, which can be initialized and read out optically. In a typical setup for magnetic sensing, the spin is first prepared in the $m_{S}=0$ state and then driven to coherent transitions (or superposition) between $m_{S}=0$ and $m_{S}=\pm 1$. An external magnetic field shifts the corresponding resonance frequencies (or the probability amplitudes between $m_{S}$’s) via the Zeeman effect. By monitoring the frequency, one can estimate the strength of the magnetic field with high precision [16, 17, 18]. This setup is a standard Ramsey interferometry for a single shot measurement [19]. Like classical sensing paradigms, $N$ independent executions of NV-center Ramsey measurements only lead to $\frac{1}{\sqrt{N}}$ error scaling due to the independent per-shot Gaussian noises. This limit is widely known as the standard quantum limit (SQL). In contrast, some recent works have demonstrated that NV-center quantum magnetometers with only correlated measurements (without using entanglement) can beat SQL and approach the Heisenberg limit (HL) with a scaling factor of $\frac{1}{N}$ [20, 21], which is the ultimate precision limit only constrained by the law of physics. Therefore, it enables highly sensitive sensing of the Earth’s magnetic field and results in accurate geo-localization.
@@ II Related Work
The foundation of any magnetic-based localization system is the magnetic anomaly map, which provides the reference field for positioning. Gupta et al. [22] established the relationship between magnetic anomaly map quality and achievable positioning accuracy via the CRLB. To implement the bound, they considered two measurement noise models: unimodal Gaussian noise and multimodal noise. Through simulations over a $250\times 400~\text{km}^{2}$ region, their results demonstrated a strong correspondence between higher magnetic gradients and lower CRLB (i.e., better accuracy). Recognizing that localization performance depends on the spatial variability of the magnetic field, Penumarti et al. [23] addressed localization uncertainty in magnetic anomaly navigation by introducing a global, uncertainty-aware path planner. Their key contribution was the generation of an entropy map from pre-existing magnetic anomaly map to statistically quantify areas of high information gain. They employed a multi-objective potential field planner that attracted the vehicle toward these low-entropy regions while simultaneously guiding it to a global goal. Hardware experiments using TurtleBot 4 equipped with a Quspin QTFM total-field magnetometer demonstrated that this method produced paths that significantly improved localization stability and reduced estimation covariance compared to existing methods.
@@
While map quality is important, accurate measurement of weak magnetic anomalies requires highly sensitive magnetometers. Bonato et al. [21] demonstrated DC magnetic field sensing with single-spin NV-center magnetometer using a real-time adaptive measurement protocol. They implemented an adaptive Ramsey interferometry protocol in which the spin readout basis was dynamically updated after each single-shot measurement using a Bayesian estimator. The protocol achieved a remarkable magnetic field sensitivity of $6.1\pm 1.7~\text{nT}/\sqrt{\text{Hz}}$ over a dynamic range of $1.78~\text{mT}$. Building on the development of high-sensitivity NV-center magnetometers, Wang et al. [12] investigated their viability as aiding sensors for INS using total magnetic intensity maps for localization. However, in practical platforms, the weak crustal magnetic signal is often corrupted by vehicle-induced interference, making signal separation a critical challenge. To improve magnetic signal fidelity, the United States Air Force Research Laboratory, in collaboration with MIT Lincoln Lab, introduced the MagNav signal enhancement challenge [24], which aimed to develop algorithms to separate the Earth’s magnetic anomaly field from strong aircraft-induced magnetic interference. A team from SandboxAQ addressed the challenge by proposing a physics-informed machine learning model that integrated the classical Tolles-Lawson (T-L) model with a Liquid Time-Constant Network (LTC) [25]. Their approach used a closed-form continuous-time (CfC) variant of the LTC to model the nonlinear dynamics of the aircraft-induced magnetic noise. Experimental results demonstrated that the proposed LTC-CfC network achieved up to a $64\%$ reduction in compensation error compared to the standard T-L model and outperformed other baseline methods.
@@
Building on these foundations, several integrated systems have demonstrated practical MagNav capabilities. Q-CTRL introduced a full-stack system [26] integrating a quantum scalar optically pumped atomic magnetometer, a classical fluxgate vector magnetometer, and a software stack for magnetic denoising and map matching. The software stack processes the magnetometer readings to perform real-time platform denoising and provides periodic position fixes by matching the cleaned magnetic signal against a known anomaly map (e.g., EMAG2v3). Across flight trials, this quantum-assured system consistently outperformed a strategic-grade INS, achieving superior bounded positioning accuracy with a best-case result of 22 meters. In ground trials, the system demonstrated a bounded positioning error approximately seven times lower than that of the INS. However, a major limitation of such map-based approaches is their dependence on pre-surveyed maps. To address this, Lee et al. [27] introduced MagSLAM, an aerial simultaneous localization and mapping framework (SLAM) that exploits the Earth’s magnetic anomaly field without relying on pre-surveyed magnetic maps. The authors adapted the FastSLAM algorithm (Rao-Blackwellized particle filter) [28] to fuse scalar magnetic anomaly measurements with an INS. Their method uses a nine-state INS error model to constrain drift, while simultaneously estimating the vehicle’s trajectory and building a local magnetic map without any prior survey. Flight tests were performed with a Cessna 208B equipped with a cesium-vapor scalar magnetometer, barometer, and high-grade INS over a $9\times 12~\text{km}$ region in Louisa, Virginia, at $150~\text{m}$ above ground level. The MagSLAM filter achieved an approximate $17~\text{m}$ distance root mean square position accuracy over a 100-minute flight.
@@ III Problem formulation
Figure 1: The CRLBs for a single-spin NV-center magnetometer and two commercial sensors (HMC5983 and AKM8975) [29], with respect to sensing time per measurement shot $\tau$. The red curve shows the NV-center CRLB after a fixed total sensing time $T_{\mathrm{total}}=1.5~\mathrm{s}$ with $T_{2}^{\ast}=1500~\mu\mathrm{s}$ and readout fidelities $F_{0}=0.88$ and $F_{1}=0.98$.
@@ IV-B Quantum Cramér–Rao Lower Bound
To quantify the gained knowledge from this measurement, we use the Fisher information. The Fisher information measures how the outcome probabilities respond to changes in the parameter of interest, $\omega$. In our case, the Fisher information, and thus the measurement precision, depends on the controllable phase $\vartheta$ of the second $\pi/2$ pulse and is given as follows [30]:
@@ IV-C Classical Magnetometer Cramér–Rao Lower Bound
To demonstrate the quantum advantage of the NV-center magnetometer, we compare its CRLB against that of established classical magnetometers. Fig. 1 shows the quantum CRLB on magnetic field sensing as a function of sensing time $\tau$ per Ramsey shot, for a fixed total sensing time $T_{total}=1.5~\mathrm{s}$. In comparison, we set the inhomogeneous dephasing time $T_{2}^{*}=1,500~\mu\mathrm{s}$, consistent with room-temperature single-spin systems [31]. The classical sensors used for this benchmark were selected from [29], which are neither dominated by external magnetic field noise (e.g., MMC246xMT) nor exhibit anisotropic resolution (e.g., YAS532, YAS529). For each device, we use the noise variance reported in [29] as the variance parameter in the classical model of Theorem 1, which yields the horizontal lines in Fig. 1. As shown in figure, the quantum CRLB is highly dependent on sensing time, achieving a minimum at an optimal value before increasing due to decoherence. The shaded regions highlight the operational windows where quantum sensor’s theoretical precision surpasses that of classical benchmarks, demonstrating quantum advantage. This trade-off motivates the need to operate the sensor in optimal sensing time.
@@ V-A System Overview
In this work, we deliberately adopt the same single-spin NV-center magnetometer setup as in our prior two-stage sensing protocol [20]. The rationale for selecting this protocol is that we have experimentally demonstrated that it can saturate the quantum CRLB as illustrated in Fig.1, thus beating SQL and approaching HL. More importantly, it achieves the best estimation accuracy among existing methods in the literature for estimating wide-range magnetic field.
@@ V-B Two-Stage Adaptive Sensing Protocol
We adopt the two-stage adaptive sensing protocol introduced in our earlier study [20] and use it as the front end of the magnetic localization pipeline. We refer readers to [20] for more details of the protocol while we only present a sketch of how it works here. The protocol comprises two phases: an offline training phase, followed by an online inference phase.
@@ VI-A Magnetic Map
In this work, we used the composite magnetic anomaly grid for the United States and Canada published by the USGS in March 2025 [32]. This dataset merges magnetic surveys [33, 34, 35] to create a magnetic anomaly map of North America with a resolution of approximately 1 km (30 arc-seconds). To extract the ROI from the broader North America dataset, we applied a geospatial clipping operation. A polygon defining the CONUS boundary was constructed from the Natural Earth Admin-1 states and provinces dataset [36]. This polygon was used to mask the original GeoTIFF, excluding Alaska, Hawaii, and Canadian territories. Fig. 3(a) shows the resulting ROI, spanning longitudes from approximately 125° to 66.5° and latitudes from 24° to 49.5°. The anomaly values within the CONUS ROI exhibit a wide dynamic range, from approximately 332.4-332.4 nT to over $430.4$ nT, reflecting the complex and varied geology of the continent.
@@ Appendix A Proof of Theorem 1: Classical Magnetometer Cramér–Rao Bound
The Fisher information $F(B)$ for a single parameter is defined as [29]: